•The 14C activity of some ancient Peruvian corn was found to be 10 disintegrations per minute per gram of C. If present-day plant life shows 15 dpm/g, how old is the Peruvian corn? The halflife of 14C is 5730 years.
1
Expert's answer
2020-06-22T11:09:09-0400
"A=A_0\cdot 2^{\frac{t}{T}}\\15=10\cdot 2^{\frac{t}{5730}}\\t=3350\ years"